# Knowledge Circuits in Pretrained Transformers

### v1

@@ Abstract
The remarkable capabilities of modern large language models are rooted in their vast repositories of knowledge encoded within their parameters, enabling them to perceive the world and engage in reasoning. The inner workings of how these models store knowledge have long been a subject of intense interest and investigation among researchers. To date, most studies have concentrated on isolated components within these models, such as the Multilayer Perceptrons and attention head. In this paper, we delve into the computation graph of the language model to uncover the knowledge circuits that are instrumental in articulating specific knowledge. The experiments, conducted with GPT2 and TinyLLAMA, has allowed us to observe how certain information heads, relation heads, and Multilayer Perceptrons collaboratively encode knowledge within the model. Moreover, we evaluate the impact of current knowledge editing techniques on these knowledge circuits, providing deeper insights into the functioning and constraints of these editing methodologies. Finally, we utilize knowledge circuits to analyze and interpret language model behaviors such as hallucinations and in-context learning. We believe the knowledge circuits hold potential for advancing our understanding of Transformers and guiding the improved design of knowledge editing [^1] Code and data are available in https://github.com/zjunlp/KnowledgeCircuits..
@@ 1 Introduction
Recently, the research community has devoted significant efforts to unraveling the knowledge storage mechanisms of these models. Various studies [12, 13, 14, 15, 16, 17, 18, 19] have been conducted to shed light on this intricate process, aiming to enhance our understanding and improve the safety and reliability of language models. The main finding in previous work is that knowledge may primarily stored in the Multilayer Perceptrons (MLPs) of Transformer-based language models. These MLPs function as a key-value neural memory, with knowledge being stored in what are termed “knowledge neurons” (KN). Based on these findings, researchers conduct Knowledge Editing [18, 20] to update the language models’ inaccurate facts, bias and unsafe content in their parametric space. Despite the initial success of these methods, there are still limitations, such as poor generalization, severe side effect, and failure to effectively utilize edited knowledge [20, 21], which motivate us to re-think previous approaches for interpreting knowledge storage in language models. Note that previous works treat the knowledge blocks as isolated components following the Restorative Theory [22], often focusing on identifying the specific blocks that stores particular knowledge. Several works [23, 24] have proposed that different types of knowledge are often located in the same areas, suggesting that the current KN thesis may be an oversimplification.
@@
To this end, instead of solely pinpointing tiny regions where the knowledge expressed can be localized, we aim to explore the cooperation between different components in Transformers like attention heads, MLPs, and embeddings, to understand how the language model stores and expresses the knowledge. Here, we introduce a new perspective: Knowledge Circuits, a critical subgraph in the language model to view the knowledge mechanism of Transformers. Note that Circuits, as a subgraph in the computation graph, has gained ever-growing attention in the mechanistic interpretability field [25]. Previous work [26, 27] has found several important circuits for specific tasks like Indirect Object Identification and Color Object. These tasks necessitate the model to search the preceding context for a matching token and copy it into the next token prediction. In this work, we aim to construct knowledge circuits that require the model to utilize stored knowledge for making predictions. Our goal is to better unveil implicit neural knowledge representations, elucidate the internal mechanisms for knowledge editing, and interpret more complex behaviors of language models. Specifically, we leverage factual recall tasks and conduct experiments across various domains, including factual, bias, linguistic, and commonsense knowledge. We utilize GPT-2 [28] and TinyLLAMA [29] to explore the potential knowledge representations and utilization mechanism in these models. As shown in Figure 1 (a), we construct knowledge circuits associated with various expressions of knowledge using the existing knowledge stored in the language model. Through those discovered knowledge circuits, we find many interesting phenomena and conclusions as follows:
@@
Knowledge circuits unveil implicit neural knowledge representations. We find that even when the knowledge circuits are used independently, the language model can recall related knowledge with a significant portion of its overall performance, demonstrating the effectiveness of those discovered knowledge representations (circuits). We also delve into specific pieces of knowledge and analyze the information flow within their respective circuits, indicating that language model tends to aggregate knowledge in the earlier to middle layers and further enhances this information in the later layers. We further uncover several special components (e.g., attention heads) in transferring information to the final token position and capturing relational information from the context (Figure 1 (b)).
@@ 3.1 Knowledge Circuits Construction
Unlike previous work [12, 18], which managed to find out the specific areas that store knowledge, we pay extra heed to the information flow that activates subsequent knowledge for answering questions. Similar to [38, 26], we write language model as a graph consisting of the input, the output, attention heads, and MLPs by considering a “residual rewrite” of the model’s computational structure. For example, this residual rewrite gives us a nearly-dense graph in GPT2-medium: one between every pair of (attention head, MLP, input, and output) nodes, except for attention heads in the same layer, which do not communicate with each other. In our paper, we concentrate on the task of answering factual open-domain questions, where the goal is to predict a target entity $o$ given a subject-relation pair $(s,r)$. A knowledge triplet $k=(s,r,o)$ is often presented to the model in the form of a natural language prompt for next token prediction (e.g., “The official language of France is ____”). The model $\mathcal{G}$ is expected to generate the target entity, which is consistent with the language model’s pretraining format. To identify the circuit that are critical for predicting the target entity $o$ for a given subject-relation pair $(s,r)$, we ablate each special edge $e_{i}=(n_{x},n_{y})$ in the computation graph $\mathcal{G}$. We then measure the impact of ablating the edge (zero ablation in our implementation) on the model’s performance using the MatchNLL loss [32] for the target $o$:
$S(e_{i})=-\log(\mathcal{G}/e_{i}(o|(s,r)))-\log(\mathcal{G}(o|(s,r)))$ (1)
@@ 3.2 Knowledge Circuits Information Analysis
Original($\mathcal{G}$) Circuit($\mathcal{C}$) Original($\mathcal{G}$) Circuit($\mathcal{C}$)
Linguistic Adj Antonym 573 0.80 1.00 $\uparrow$ 0.00 0.40 $\uparrow$
world first letter 432 1.00 0.88 0.36 0.16
world last letter 230 1.00 0.72 0.76 0.76
Commonsense object superclass 102 1.00 0.68 0.64 0.52
fruit inside color 433 1.00 0.20 0.93 0.13
work location 422 1.00 0.70 0.10 0.10
Factual Capital City 451 1.00 1.00 0.00 0.00
Landmark country 278 1.00 0.60 0.16 0.36 $\uparrow$
Country Language 329 1.00 1.00 0.16 0.75 $\uparrow$
Person Native Language 92 1.00 0.76 0.50 0.76 $\uparrow$
Bias name religion 423 1.00 0.50 0.42 0.42
occupation age 413 1.00 1.00 1.00 1.00
occupation gender 226 1.00 0.66 1.00 0.66
name birthplace 276 1.00 0.57 0.07 0.57 $\uparrow$
Avg 0.98 0.73 0.44 0.47 $\uparrow$
@@ 3.3 Knowledge Circuits Experimental Settings
We conduct experiments on GPT-style models, including GPT-2 medium and large. We also conduct primary experiments on TinyLLaMA [29] to validate the effectiveness of different architectures. We utilize the Automated Circuit Discovery [32] toolkit to build a circuit as an initiative of our analysis and leverage transformer lens [41] to further analyze the results. Specially, we simply employ the MatchNLL [32] as the metric to detect the effect of the given node and edge and use zero ablation to knock out the specific computation node in the model’s computation graph.
@@
A discovered knowledge circuit is deemed an accurate representation of a specific area within the transformer’s knowledge storage, thus, it should be capable of representing the knowledge independently. Following [32], we leverage the completeness of a circuit which refers to its ability to independently reproduce the behavior or predictions of the full model for the relevant tasks. This property is assessed by examining whether the identified subgraph corresponds to the underlying algorithm implemented by the neural network. To evaluate completeness, we first construct the circuit using the validation data $D_{val}$ for a specific knowledge type and then test its performance on the test split $D_{test}$ in isolation. By doing so, we can observe any changes in performance compared to the original model. We use the Hit@10 metric to measure the rank of the target entity $o$ among the top 10 predicted tokens:
@@
In this work, we focus on the knowledge that already stored in the language model. We utilize the dataset provided by LRE [42] and consider different kinds of knowledge, including linguistic, commonsense, fact, and bias. We evaluate whether the knowledge is present in the language model’s parameters under zero-shot settings using the Hit@10 metric to sample knowledge from the validation set, which is used to construct the knowledge circuit. The data statistics are in Appendix A.
@@ 4 Knowledge Circuits Unveil Implicit Neural Knowledge Representations
We report the results of GPT2-Medium in Table 1, wich indicates that with only less than 10% of the original knowledge circuit’s subgraph, the model can maintain over 70% of its original performance. One of the most fascinating observations is the performance improvement seen on several test datasets. For instance, the Landmark-country relation metric increases from 0.16 to 0.36. This suggests that the discovered knowledge circuits may encapsulate the relevant knowledge, and the model’s performance on these tasks could have been hindered by noise from other components. We proceed to analyze the layer distribution of the original model $\mathcal{G}$ to understand the average percentage of nodes that are activated within the circuit for different knowledge domains. From Figure 2, we observe that attention and MLPs are more active in the lower layers of the network, where the language model processes the input and extracts general information. To gain a more comprehensive view of the information processing, we compute the average $\operatorname{rank}_{o}$ change of the target token in the $D_{val}$ across the layers and report the results in Figure 6. This analysis reveals the phenomenon of early decoding [40], suggesting that by the middle to the latest layers, the target entity is already present in the residual stream, and the subsequent layers in the Transformer are designed to increase the probability of the current token (See discussion in the running example).
@@
From the discovered knowledge circuits, we can find several important attention heads that demonstrate specific behavior, including the mover head [31], relation head [17, 43] and mixture head [17, 43] (Definitions in Appendix B.2). We think that these components would be accumulated by the MLP in the model and discuss the behavior of these special heads in the running example part. We list some of these special components in Table 3 in Appendix. The different attention heads are responsible for expressing specific types of knowledge and may be activated by different facts. In our experiments with GPT-2 Medium and GPT-2 Large, we find that knowledge is distributed across several layers’ attention heads and MLP matrices, suggesting that the target knowledge appears to have been accumulated throughout the GPT-2 model. Conversely, in TinyLLAMA, the special components are more concentrated. As depicted in Figure 6, the rank of the target entity in TinyLLAMA experiences a sharp decline around several layers, whereas in the GPT2 model, the decline is more gradual. We hypothesize that this discrepancy may be attributed to the model’s knowledge capacity [44] and warrants further investigation.
@@
Figure 3: The rank and probability of the target entity $o$ when unembedding the intermediate layer’s output for the fact “The official language of France is French”. For a more comprehensive view, we additionally plot the probability of the subject entity at the last token position and the target entity’s rank at the subject position.
We present a case and analyze the specific behaviors of components within the identified knowledge circuits. Taking the factual knowledge “The official language of France is French” as an example, we visualize the knowledge circuit in Figure 1. To express the information flow within the model more effectively, we have plotted the rank and probability of the target entity $o$ at each layer when it is mapped into the embedding space, in Figure 3. After MLP 17, the target knowledge emerges as the top token in the residual stream and after that layer, it undergoes an increased probability. The edges connected to MLP17 are (L14H13 $\rightarrow$ MLP17), (L14H7 $\rightarrow$ MLP17), and (L15H0 $\rightarrow$ MLP17) . Here, the L14H13 is a relation head that focuses on the relation token in the context. The output of this head is relation-related tokens such as “language” and “Language”. The attention head L14H7 is a mover head that moves the information from the subject position “France” to the last token. Previous work [31, 19] has introduced this mover head as an argument parser, which moves “France” to the last token, and the subsquent MLP conducts a function application to map “France” to “French”. An intriguing observation is that we can find the output of this head already contains the target entity, which significantly contributes to the final output (L14H7 $\rightarrow$ Output). Also, we see the probability of the subject token at the last token is nearly zero across these layers. Hence, instead of the argument parser, we consider this mover head as an extract head proposed by Geva et al. [13], which aims to extract the related-information from the subject token’s position. In the subsequent knowledge editing experiments, we can observe changes in the behavior of these types of heads. Instead of extraction in the later layers proposed by Geva et al. [13], we notice a gradual decrease in rank across all early-to-middle layers. The MLP17 combines information from previous tokens and integrates this information to prioritize the target token at the top rank.

### v2

@@ Abstract
The remarkable capabilities of modern large language models are rooted in their vast repositories of knowledge encoded within their parameters, enabling them to perceive the world and engage in reasoning. The inner workings of how these models store knowledge have long been a subject of intense interest and investigation among researchers. To date, most studies have concentrated on isolated components within these models, such as the Multilayer Perceptrons and attention head. In this paper, we delve into the computation graph of the language model to uncover the knowledge circuits that are instrumental in articulating specific knowledge. The experiments, conducted with GPT2 and TinyLLAMA, have allowed us to observe how certain information heads, relation heads, and Multilayer Perceptrons collaboratively encode knowledge within the model. Moreover, we evaluate the impact of current knowledge editing techniques on these knowledge circuits, providing deeper insights into the functioning and constraints of these editing methodologies. Finally, we utilize knowledge circuits to analyze and interpret language model behaviors such as hallucinations and in-context learning. We believe the knowledge circuits hold potential for advancing our understanding of Transformers and guiding the improved design of knowledge editing [^1] Code and data are available in https://github.com/zjunlp/KnowledgeCircuits..
@@ 1 Introduction
Recently, the research community has devoted significant efforts to unraveling the knowledge storage mechanisms of these models. Various studies [12, 13, 14, 15, 16, 17, 18, 19] have been conducted to shed light on this intricate process, aiming to enhance our understanding and improve the safety and reliability of language models. The main finding in previous work is that knowledge may primarily stored in the Multilayer Perceptrons (MLPs) of Transformer-based language models. These MLPs function as a key-value neural memory, with knowledge being stored in what are termed “knowledge neurons” (KN). Based on these findings, researchers conduct Knowledge Editing [18, 20] to update the language models’ inaccurate facts, bias and unsafe content in their parametric space. Despite the initial success of these methods, there are still limitations, such as poor generalization, severe side effects, and failure to effectively utilize edited knowledge [20, 21], which motivate us to re-think previous approaches for interpreting knowledge storage in language models. Note that previous works treat the knowledge blocks as isolated components following the Restorative Theory [22], often focusing on identifying the specific blocks that store particular knowledge. Several works [23, 24] have proposed that different types of knowledge are often located in the same areas, suggesting that the current KN thesis may be an oversimplification.
@@
To this end, instead of solely pinpointing tiny regions where the knowledge expressed can be localized, we aim to explore the cooperation between different components in Transformers like attention heads, MLPs, and embeddings, to understand how the language model stores and expresses the knowledge. Here, we introduce a new perspective: Knowledge Circuits, a critical subgraph in the language model to view the knowledge mechanism of Transformers. Note that Circuit, as a subgraph in the computation graph, has gained ever-growing attention in the mechanistic interpretability field [25]. Previous work [26, 27] has found several important circuits for specific tasks like Indirect Object Identification and Color Object Identification. These tasks necessitate the model to search the preceding context for a matching token and copy it into the next token prediction. In this work, we aim to construct knowledge circuits that require the model to utilize stored knowledge for making predictions. Our goal is to better unveil implicit neural knowledge representations, elucidate the internal mechanisms for knowledge editing, and interpret more complex behaviors of language models. Specifically, we leverage factual recall tasks and conduct experiments across various domains, including factual, social bias, linguistic, and commonsense knowledge. We utilize GPT-2 [28] and TinyLLAMA [29] to explore the potential knowledge representations and utilization mechanisms in these models. As shown in Figure 1 (a), we construct knowledge circuits associated with various expressions of knowledge using the existing knowledge stored in the language model. Through those discovered knowledge circuits, we find many interesting phenomena and conclusions as follows:
@@
Knowledge circuits unveil implicit neural knowledge representations. We find that even when the knowledge circuits are used independently, the language model can recall related knowledge with a significant portion of its overall performance, demonstrating the effectiveness of those discovered knowledge representations (circuits). We also delve into specific pieces of knowledge and analyze the information flow within their respective circuits, indicating that the language model tends to aggregate knowledge in the earlier to middle layers and further enhances this information in the later layers. We further uncover several special components (e.g., mover heads and relation heads) in transferring information to the final token position and capturing relational information from the context (Figure 1 (b)).
@@ 3.1 Knowledge Circuits Construction
Unlike previous work [12, 18], which managed to find out the specific areas that store knowledge, we pay extra heed to the information flow that activates subsequent knowledge for answering questions. Similar to [38, 26], we write language model as a graph consisting of the input, the output, attention heads, and MLPs by considering a “residual rewrite” of the model’s computational structure. For example, this residual rewrite gives us a nearly-dense graph in GPT2-medium: one between every pair of (attention head, MLP, input, and output) nodes, except for attention heads in the same layer, which do not communicate with each other. In our paper, we concentrate on the task of answering factual open-domain questions, where the goal is to predict a target entity $o$ given a subject-relation pair $(s,r)$. A knowledge triplet $k=(s,r,o)$ is often presented to the model in the form of a natural language prompt for next token prediction (e.g., “The official language of France is ____”). The model $\mathcal{G}$ is expected to generate the target entity, which is consistent with the language model’s pretraining format. To identify the circuit that is critical for predicting the target entity $o$ for a given subject-relation pair $(s,r)$, we ablate each special edge $e_{i}=(n_{x},n_{y})$ in the computation graph $\mathcal{G}$. We then measure the impact of ablating the edge (zero ablation in our implementation) on the model’s performance using the MatchNLL loss [32] for the target $o$:
$S(e_{i})=\log(\mathcal{G}(o|(s,r)))-\log(\mathcal{G}/e_{i}(o|(s,r)))$ (1)
@@ 3.2 Knowledge Circuits Information Analysis
Original($\mathcal{G}$) Circuit($\mathcal{C}$) Original($\mathcal{G}$) Random Circuit($\mathcal{C}$)
Linguistic Adj Antonym 573 0.80 1.00 $\uparrow$ 0.00 0.00 0.40 $\uparrow$
word first letter 432 1.00 0.88 0.36 0.00 0.16
word last letter 230 1.00 0.72 0.76 0.00 0.76
Commonsense object superclass 102 1.00 0.68 0.64 0.00 0.52
fruit inside color 433 1.00 0.20 0.93 0.00 0.13
work location 422 1.00 0.70 0.10 0.00 0.10
Factual Capital City 451 1.00 1.00 0.00 0.00 0.00
Landmark country 278 1.00 0.60 0.16 0.00 0.36 $\uparrow$
Country Language 329 1.00 1.00 0.16 0.00 0.75 $\uparrow$
Person Native Language 92 1.00 0.76 0.50 0.00 0.76 $\uparrow$
Bias name religion 423 1.00 0.50 0.42 0.00 0.42
occupation age 413 1.00 1.00 1.00 0.00 1.00
occupation gender 226 1.00 0.66 1.00 0.00 0.66
name birthplace 276 1.00 0.57 0.07 0.00 0.57 $\uparrow$
Avg 0.98 0.73 0.44 0.00 0.47 $\uparrow$
@@ 3.3 Knowledge Circuits Experimental Settings
We conduct experiments on GPT-style models, including GPT-2 medium and large. We also conduct primary experiments on TinyLLaMA [29] to validate the effectiveness of different architectures. We utilize the Automated Circuit Discovery [32] toolkit to build a circuit as an initiative of our analysis and leverage transformer lens [41] to further analyze the results. Specifically, we simply employ the MatchNLL [32] as the metric to detect the effect of the given node and edge and use zero ablation to knock out the specific computation node in the model’s computation graph.
@@
A discovered knowledge circuit is deemed an accurate representation of a specific area within the transformer’s knowledge storage, thus, it should be capable of representing the knowledge independently. Following [32], we leverage the completeness of a circuit, which refers to its ability to independently reproduce the behavior or predictions of the full model for the relevant tasks. This property is assessed by examining whether the identified subgraph corresponds to the underlying algorithm implemented by the neural network. To evaluate completeness, we first construct the circuit using the validation data $D_{val}$ for a specific knowledge type and then test its performance on the test split $D_{test}$ in isolation. By doing so, we can observe any changes in performance compared to the original model. We use the Hit@10 metric to measure the rank of the target entity $o$ among the top 10 predicted tokens:
@@
In this work, we focus on the knowledge that is already stored in the language model. We utilize the dataset provided by LRE [42] and consider different kinds of knowledge, including linguistic, commonsense, fact, and bias. We evaluate whether the knowledge is present in the language model’s parameters under zero-shot settings using the Hit@10 metric to sample knowledge from the validation set, which is used to construct the knowledge circuit. The data statistics are in Appendix A.
@@ 4 Knowledge Circuits Unveil Implicit Neural Knowledge Representations
We report the results of GPT2-Medium in Table 1, which indicates that with only less than 10% of the original knowledge circuit’s subgraph, the model can maintain over 70% of its original performance. Additionally, we compute the random circuits by randomly deciding whether the edge should be removed and making sure the graph is connected. The random circuit is the same size as the circuit we discovered using our method. From the table, we can see that the random circuit failed to maintain the model’s performance, which further enhanced the robustness and efficacy of our methods. One of the most fascinating observations is the performance improvement seen on several test datasets. For instance, the Landmark-country relation metric increases from 0.16 to 0.36. This suggests that the discovered knowledge circuits may encapsulate the relevant knowledge, and the model’s performance on these tasks could have been hindered by noise from other components. We proceed to analyze the layer distribution of the original model $\mathcal{G}$ to understand the average percentage of nodes that are activated within the circuit for different knowledge domains. From Figure 2, we observe that attention and MLPs are more active in the lower layers of the network, where the language model processes the input and extracts general information. To gain a more comprehensive view of the information processing, we compute the average $\operatorname{rank}_{o}$ change of the target token in the $D_{val}$ across the layers and report the results in Figure 7. This analysis reveals the phenomenon of early decoding [40], suggesting that by the middle to the latest layers, the target entity is already present in the residual stream, and the subsequent layers in the Transformer are designed to increase the probability of the current token (See discussion in the running example).
@@
From the discovered knowledge circuits, we can find several important attention heads that demonstrate specific behavior, including the mover head [31], relation head [17, 43] and mixture head [17, 43] (more definitions in Appendix B.2). Mover Head [31, 27] focuses on the last token of the context and attends to the subject token, functioning as a mover to transfer information, while Relation Head [17] attends to the relation token in the context and produces some relation-related tokens that would guide the behavior of the following components. We think that these components would be accumulated by the MLP in the model, and the behavior of these special heads will be discussed in the running example part. We list some of these special components in Table 4 in Appendix. The different attention heads are responsible for expressing specific types of knowledge and may be activated by different facts. In our experiments with GPT-2 Medium and GPT-2 Large, we find that knowledge is distributed across several layers’ attention heads and MLP matrices, suggesting that the target knowledge appears to have been accumulated throughout the GPT-2 model. Conversely, in TinyLLAMA, the special components are more concentrated. As depicted in Figure 7, the rank of the target entity in TinyLLAMA experiences a sharp decline around several layers, whereas in the GPT2 model, the decline is more gradual. We hypothesize that this discrepancy may be attributed to the model’s knowledge capacity [44] and warrants further investigation.
@@
We present a case and analyze the specific behaviors of components within the identified knowledge circuits. In particular, we find some special attention heads in the model such as the mover head and the relation head. We demonstrate the function of these heads in Figure 6 by ablating them from the circuit. Taking the factual knowledge “The official language of France is French” as an example, we visualize the knowledge circuit in Figure 1. To express the information flow within the model more effectively, we have plotted the rank and probability of the target entity $o$ at each layer when it is mapped into the embedding space, in Figure 3. From this figure, we can see that after MLP 17, the target knowledge emerges as the top token in the residual stream, and after that layer, it undergoes an increased probability. The edges connected to MLP17 are (L14H13 $\rightarrow$ MLP17), (L14H7 $\rightarrow$ MLP17), and (L15H0 $\rightarrow$ MLP17) . Here, the L14H13 is a relation head that focuses on the relation token in the context. The output of this head is relation-related tokens such as “language” and “Language”. The attention head L14H7 is a mover head that moves the information from the subject position “France” to the last token. Previous work [31, 19] has introduced this mover head as an argument parser, which moves “France” to the last token, and the subsequent MLP conducts a function application to map “France” to “French”. An intriguing observation is that we can find the output of this head already contains the target answer entity, which significantly contributes to the final output (L14H7 $\rightarrow$ Output). Also, we see the probability of the subject token in Figure 3 at the last token is nearly zero across these layers. Hence, instead of the argument parser function, we consider this mover head as an extract head proposed by Geva et al. [13], which aims to extract the related-information from the subject token’s position. In the subsequent knowledge editing experiments, we can observe changes in the behavior of these types of heads.
Figure 3: The rank and probability of the target entity $o$ at both the last subject token and the last token position when unembedding the intermediate layer’s output for the fact “The official language of France is French”.
Additionally, instead of extraction in the later layers proposed by Geva et al. [13], we notice a gradual decrease in rank across all early-to-middle layers. The MLP17 combines information from previous tokens and integrates this information to prioritize the target token at the top rank.